# Revolutionizing Future Connectivity: A Contemporary Survey on AI-empowered Satellite-based Non-Terrestrial Networks in 6G

### v3

@@ V-B3 Security - Physical Layer Authentication
Due to the new interfaces introduced by satellite-integrated terrestrial architectures, various spoofing and replay attacks can be launched using these interfaces. Spoofing attacks involve an attacker satellite impersonating a legitimate one, while replay attacks involve the retransmission of previously intercepted messages to deceive users. Generally, in terrestrial networks, these kinds of attacks are detected and mitigated by using standard cryptographic techniques, a concept also investigated in satellite communications [169, 170]. However, when it comes to NTN-integrated future 6G networks, these conventional cryptographic methods face several challenges. Firstly, these techniques are computationally intensive and, thereby challenging to implement in satellites due to their limited onboarding capabilities. Secondly, the highly dynamic and massive scale network topology of NTNs, particularly for enabling IoT devices, necessitates significant modifications in network protocol design and introduces overheads that may not be practical to manage with existing architectures Also, these cryptographic techniques often assume that attackers lack the computational resources to break the encryption. However, with ongoing advances in quantum computing research, these assumptions may no longer hold in the future, presenting yet another challenge that needs to be addressed.
@@
Physical layer authentication offers a promising alternative to these conventional techniques. In [171], Wyner introduced the concept of physical layer authentication where the message was encoded in such a way that the mutual information between the legitimate channel and wiretap channel is maximized. This encoding generally captures the unique characteristics of the channel between the user and the legitimate transmitter, serving as a means to verify the transmitter’s identity. This technique has been already explored in terrestrial networks using the CSI information as radio signatures for the transmitter devices [172, 173, 174]. However, the prevalence of Line of Sight (LoS) paths in satellite networks makes radio fingerprinting using channel fading information from CSI impractical. Furthermore, due to the high mobility of NGSO satellites, as discussed in Section II-E, the high Doppler Shift is introduced in received signals, which can be used to verify the identity of the legitimate satellites. In [175], a maximum likelihood estimation and uniform quantizer are used to obtain the secret key bits from the Doppler frequency shifts, which is used in the authentication of legitimate satellites [176] In [177], an orbital information – time difference of arrival-based authentication mechanism is introduced providing low false authentication rates.
@@
In the recent past, various DL techniques have shown lofty promises in the field of extracting features from noisy data, which is also leveraged in this field using different ML models. In [178], CNNs and Autoencoders are used to extract the necessary channel features for physical layer authentication of legitimate satellites. In [179, 180, 181], both the received signal power and Doppler shift are used for radio fingerprinting using SVMs providing improved authentication rates.
@@ V-B4 Security - Intrusion Detection
To address these challenges, DL techniques are explored to efficiently detect security threats using various innovative approaches. In the study presented in [182], an ensemble model combining Random Forest (RF) and MLP is developed to improve the performance of security threat detection across diverse datasets for satellite communications. [183] leverages critical feature selection driven by RF to streamline complexity and enhance the relevance of features before the detection phase. These features are then forwarded to different NN architectures, including LSTM, GRU, RF, and ANN enabling robust security threat detection. These models are tested on different datasets where GRU-empowered threat detection models exhibit superior performances by capturing temporal behavioral patterns. In [184], a UL approach using LSTM networks is explored, which can not only detect unforeseen security threats but also does not need any labeled data. In another study as shown in [185], two SL and five UL approaches are considered for threat detection to show the effectiveness of ML techniques. In [186], a DDPG-based DRL framework is considered where the agents decide whether the aerial platform is malicious or not (actions) based on their behavior (states) and the system condition (rewards) for threat detection. Recognizing the computational constraints of satellites and Internet of Things (IoT) devices, federated learning approaches are also investigated as detailed in [187, 188, 189] for threat detection.
@@ V-B5 Security - Anti-jamming
Satellites are vulnerable to jamming threats due to their predictable and periodic visibility in NTNs, so anti-jamming approaches are important to tackle this challenge. Conventional spread spectrum techniques are used in anti-jamming for satellite networks. However, they are not very useful in dealing with new smart jamming attacks which can adjust their actions based on the network feedback. Various RL techniques are adopted to tackle these problems in an efficient manner. In [190, 191], a hierarchical anti-jamming Stackelberg game is introduced for routing anti-jamming problems which is later solved by providing fast anti-jamming decisions using a DRL-based routing algorithm for satellites. in [192], a DL-based jamming detection algorithm is proposed for satellite navigation systems. In [193], an anti-jamming coalition game is formed to decrease energy consumption, and suboptimal jamming policies are obtained by RL approaches. In [194], ML-aided cognitive anti-jamming communication is designed, developed, and tested on real satellite-ground links.
@@ V-C1 Beam Hopping
Modern communication satellites form multiple beams to support a large number of users over a large area through spatial multiplexing into different NTN cells. Each satellite can effectively reuse the allocated spectrum with very low co-channel interference as well as provide strong signals at the ground user terminals with relatively low transmission power using beamforming techniques. However, due to the high cost and low availability of onboard processing computing resources in satellite systems, mostly simple fixed beam allocation policies are used in traditional satellite communication. These strategies lack the flexibility to adapt to the temporal and spatial variation of traffic demands in real satellite networks. Beam hopping is a technique for allocating beams in a flexible manner so that these changes can be addressed efficiently. It refers to a procedure for activating different beams according to the current demands of an NTN cell covered by those beams, so effectively hopping the set of active beams from one combination to another [195]. In Figure 13, a simple beam-hopping scenario is depicted, where we have different NTN cells with varying demands. We classify the cells into three different categories, e.g., high, medium, and low, based on their traffic demands. In the first scenario, the low-demand NTN cells, e.g. cell , have less number of active beams than high-demand NTN cells, e.g. cell , even lesser than moderate-demand NTN cells, e,g, cell or cell . However, due to mobility or change in traffic patterns, the traffic demand in cell reduces and in cell increases. As a result, we can see the intensity of the beams also changes accordingly in these two cells at a later time, and a new beam-hopping pattern emerges.
@@
Throughput |
Tx Delay |
Delay Fairness |
Capacity-Demand Ratio |
SL | UL | RL |
RL Model |
DL Model |
[110] |
FCNN |
[111] |
FCNN |
[196] |
DQN |
FCNN |
[197] |
DQN |
CNN |
[198] |
Proximal Policy Optimization |
FCNN |
[199] |
DQN |
CNN |
[200] |
DQN |
CNN |
[201] |
DDQN |
FCNN |
The key question of beam hopping is to find out which beams need to be activated when and for how long while maximizing the network performances given the capacity constraints [202]. This can be effectively formulated as an optimization problem considering different network performance metrics such as system throughput, delay, fairness, etc. as the objective(s) along with power and spectrum constraints. In [203], a convex optimization framework with an objective to match the system capacity to traffic demand along with power allocation constraints is considered. This yields a close-form solution giving insights into resource allocation policies for maximizing network performances from different perspectives. However, from the perspective of real networks, the convex objective function is not very realistic, so the results are not applicable to real networks in a straightforward manner. Assuming the non-convexity of the problem, obtaining a globally optimal solution with efficient algorithms gets difficult. In [109], the steepest gradient descent algorithm is chosen to get the sub-optimal solution using the optimal set of precoding vectors. Some heuristic iterative approaches are also proposed in [108], [204], [205] to tackle these non-convex problems in a practical and feasible manner. Different meta-heuristic approaches like Genetic Algorithm (GA) [206], Simulated Annealing (SA) Algorithm [207], Particle Swarm Optimization (PSO) [208], and combined metaheuristic approaches like GA-SA [209] have been considered to generate suboptimal solutions with a reduced amount of computational complexity.
@@
In [110], [111], an SL approach is considered by forming labeled datasets with beam hopping patterns as outputs and channel matrix, transmission power, and traffic demand as inputs. First, a mixed integer linear problem formulation for matching the offered capacity to traffic demands is reduced to a simple linear programming problem. A training dataset is generated using conventional optimization algorithms and a DL model is trained on this dataset by considering beam hopping patterns as labels. Furthermore, the optimization framework can be potentially transformed into an RL problem to capture the optimal beam-hopping pattern in a learning environment. In [196],[197] the transmission delay is minimized considering the power and beam allocation constraints using a DRL approach. The state space consists of the average transmission delay and the buffer length with beam hopping pattern as actions and the negative Hadamard product of the current states, the negative of total queuing delay as the reward function. In [198], a combined DRL-metaheuristic approach is considered to optimize both the throughput and delay fairness while at the same time designing different reward functions for the two cases. In [199], [200], a network consisting of real-time and non-real-time traffic is considered. A multi-objective problem minimizing the transmission delay for real-time traffics, maximizing the throughput for non-real-time traffics as well as overall delay fairness is considered. Individual reward functions are designed to capture each of the goals. In [201], a cooperative multi-agent framework is considered to dynamically allocate the power and bandwidth to illuminating beams optimizing throughput and delay fairness using a DDQN. In table V, we summarize the AI approaches for beam-hopping in NTN. As traffic demand changes with time, recursive architectures such as RNN, ESN, etc. should be also explored to design the NN for DL architectures used to address beam-hopping issues. Also, distributed learning architectures can be useful to design efficient beam-hopping schemes.
@@ V-C2 Spectrum Sharing
Spectrum Sensing |
Spectrum Occupancy Prediction |
Spectrum Access |
SL | UL | RL |
RL Model |
DL Model |
[210] |
CNN-BiLSTM |
[211] |
CNN |
[212] |
CNN-LSTM |
[213] |
CNN-LSTM |
[214] |
Modified Q-Learning |
[215] |
SVM-CNN |
[216] |
MADDPG |
FCNN |
@@
In TNTN, the spectrum-sharing phenomenon needs more attention because we have a hierarchical network scenario consisting of non-terrestrial and terrestrial BSs as shown in Figure 3. To support this complex topology in a single framework, we need to come up with efficient spectrum-sharing strategies causing low interference to the users [217]. In conventional spectrum sharing methods, we use efficient frequency reuse, leveraging directional antennas, adaptive power control, etc. methods to mitigate the effect of CCI. However, traditional four-color frequency reuse can effectively reduce the level of interference at the expense of more spectrum. beamforming approach can reduce the interference greatly, but that too comes at the cost of increasing complexity.
@@
To tackle this situation, a process called spectrum sensing is introduced in cognitive radio networks, where the unlicensed users can sense the occupancy status of the target band using some radio sensing method [218]. The popular spectrum sensing methods are Energy Detection (ED)[219], Cyclo-Stationary Detection (CSD) [220], Eigen Value-based Detection (EVD) [221] etc. However, these methods either are simple with poor performance in low SNR scenarios (ED) or provide good performance but with more computational complexity (CSD and EVD). For these reasons, ML has been adopted for spectrum sharing to capture the correlation with a reduced computational complexity which can be extended to integrated satellite-terrestrial network scenarios [217].
@@
Different intelligent learning approaches are adopted to tackle the spectrum-sharing problem for next-generation TNTN networks [222]. In [210], a spectrum-sharing strategy is developed for LEO satellites from the GEO satellite spectrum historical occupancy data using a CNN-BiLSTM model. Here the LEO satellite users are considered as unlicensed secondary users and the GEO satellite users are considered as the licensed primary users. In [212], a CNN-LSTM-based spectrum sensing method is introduced for satellites to capture the spatial and temporal correlation effectively for spectrum occupancy of satellite systems. In [213], a CNN-LSTM model is introduced to predict the frequency assignment for satellites based on historical data. In [214], a modified Q-Learning algorithm is used in an RL setup for the adaptive selection of access and modulation schemes for NGSO satellites in an NGSO-GEO system. In [215], an SVM model is first used for low complexity spectrum sensing, then a CNN-based spectrum prediction model based on historical data is developed. In [216], a cooperative MADRL framework is considered for bandwidth management in a game-theoretic model minimizing inter-beam interference. In [211], a CNN-based spectrum reconstruction method from incomplete data is discussed for satellite networks. In table VI, we summarize the AI approaches for spectrum sharing in TNTN. However, these spectrum sensing decisions need to be in real-time to increase the overall throughput of the secondary users; this means the conventional LSTM architectures need to be replaced by efficient low-complexity ESN architectures to tackle this in an online manner. Furthermore, spatial spectrum sharing scenarios need to be also considered along with the state of art temporal spectrum sharing scenarios leveraging the benefits of future 3D SAGIN networks.

### v4

@@ V-B3 Security - Physical Layer Authentication
Due to the new interfaces introduced by satellite-integrated terrestrial architectures, various spoofing and replay attacks can be launched using these interfaces. Spoofing attacks involve an attacker satellite impersonating a legitimate one, while replay attacks involve the retransmission of previously intercepted messages to deceive users. Generally, in terrestrial networks, these kinds of attacks are detected and mitigated by using standard cryptographic techniques, a concept also investigated in satellite communications [ntn-pla-gen1, ntn-pla-gen2]. However, when it comes to NTN-integrated future 6G networks, these conventional cryptographic methods face several challenges. Firstly, these techniques are computationally intensive and, thereby challenging to implement in satellites due to their limited onboarding capabilities. Secondly, the highly dynamic and massive scale network topology of NTNs, particularly for enabling IoT devices, necessitates significant modifications in network protocol design and introduces overheads that may not be practical to manage with existing architectures Also, these cryptographic techniques often assume that attackers lack the computational resources to break the encryption. However, with ongoing advances in quantum computing research, these assumptions may no longer hold in the future, presenting yet another challenge that needs to be addressed.
@@
Physical layer authentication offers a promising alternative to these conventional techniques. In [pla-basic], Wyner introduced the concept of physical layer authentication where the message was encoded in such a way that the mutual information between the legitimate channel and wiretap channel is maximized. This encoding generally captures the unique characteristics of the channel between the user and the legitimate transmitter, serving as a means to verify the transmitter’s identity. This technique has been already explored in terrestrial networks using the CSI information as radio signatures for the transmitter devices [pla-ter1, pla-ter2, pla-ter3]. However, the prevalence of Line of Sight (LoS) paths in satellite networks makes radio fingerprinting using channel fading information from CSI impractical. Furthermore, due to the high mobility of NGSO satellites, as discussed in Section II-E, the high Doppler Shift is introduced in received signals, which can be used to verify the identity of the legitimate satellites. In [ntn-pla-nml1], a maximum likelihood estimation and uniform quantizer are used to obtain the secret key bits from the Doppler frequency shifts, which is used in the authentication of legitimate satellites [ntn-pla-nml2] In [ntn-pla-orbit], an orbital information – time difference of arrival-based authentication mechanism is introduced providing low false authentication rates.
@@
In the recent past, various DL techniques have shown lofty promises in the field of extracting features from noisy data, which is also leveraged in this field using different ML models. In [ntn-pla2], CNNs and Autoencoders are used to extract the necessary channel features for physical layer authentication of legitimate satellites. In [ntn-pla1, ntn-pla3, ntn-pla4], both the received signal power and Doppler shift are used for radio fingerprinting using SVMs providing improved authentication rates.
@@ V-B4 Security - Intrusion Detection
To address these challenges, DL techniques are explored to efficiently detect security threats using various innovative approaches. In the study presented in [ntn-id-rfmlp], an ensemble model combining Random Forest (RF) and MLP is developed to improve the performance of security threat detection across diverse datasets for satellite communications. [ntn-id-rfdl] leverages critical feature selection driven by RF to streamline complexity and enhance the relevance of features before the detection phase. These features are then forwarded to different NN architectures, including LSTM, GRU, RF, and ANN enabling robust security threat detection. These models are tested on different datasets where GRU-empowered threat detection models exhibit superior performances by capturing temporal behavioral patterns. In [ntn-id-lstm-ul], a UL approach using LSTM networks is explored, which can not only detect unforeseen security threats but also does not need any labeled data. In another study as shown in [ntn-id-lstm-rnn], two SL and five UL approaches are considered for threat detection to show the effectiveness of ML techniques. In [ntn-id-drl], a DDPG-based DRL framework is considered where the agents decide whether the aerial platform is malicious or not (actions) based on their behavior (states) and the system condition (rewards) for threat detection. Recognizing the computational constraints of satellites and Internet of Things (IoT) devices, federated learning approaches are also investigated as detailed in [ntn-id-fed1, ntn-id-fed2, ntn-id-fed3] for threat detection.
@@ V-B5 Security - Anti-jamming
Satellites are vulnerable to jamming threats due to their predictable and periodic visibility in NTNs, so anti-jamming approaches are important to tackle this challenge. Conventional spread spectrum techniques are used in anti-jamming for satellite networks. However, they are not very useful in dealing with new smart jamming attacks which can adjust their actions based on the network feedback. Various RL techniques are adopted to tackle these problems in an efficient manner. In [ntn-aj1, ntn-aj3], a hierarchical anti-jamming Stackelberg game is introduced for routing anti-jamming problems which is later solved by providing fast anti-jamming decisions using a DRL-based routing algorithm for satellites. in [ntn-aj4], a DL-based jamming detection algorithm is proposed for satellite navigation systems. In [ntn-aj2], an anti-jamming coalition game is formed to decrease energy consumption, and suboptimal jamming policies are obtained by RL approaches. In [ntn-aj5], ML-aided cognitive anti-jamming communication is designed, developed, and tested on real satellite-ground links.
@@ V-C1 Beam Hopping
Modern communication satellites form multiple beams to support a large number of users over a large area through spatial multiplexing into different NTN cells. Each satellite can effectively reuse the allocated spectrum with very low co-channel interference as well as provide strong signals at the ground user terminals with relatively low transmission power using beamforming techniques. However, due to the high cost and low availability of onboard processing computing resources in satellite systems, mostly simple fixed beam allocation policies are used in traditional satellite communication. These strategies lack the flexibility to adapt to the temporal and spatial variation of traffic demands in real satellite networks. Beam hopping is a technique for allocating beams in a flexible manner so that these changes can be addressed efficiently. It refers to a procedure for activating different beams according to the current demands of an NTN cell covered by those beams, so effectively hopping the set of active beams from one combination to another [bh-ov]. In Figure 13, a simple beam-hopping scenario is depicted, where we have different NTN cells with varying demands. We classify the cells into three different categories, e.g., high, medium, and low, based on their traffic demands. In the first scenario, the low-demand NTN cells, e.g. cell , have less number of active beams than high-demand NTN cells, e.g. cell , even lesser than moderate-demand NTN cells, e,g, cell or cell . However, due to mobility or change in traffic patterns, the traffic demand in cell reduces and in cell increases. As a result, we can see the intensity of the beams also changes accordingly in these two cells at a later time, and a new beam-hopping pattern emerges.
@@
| Throughput | Tx Delay | Delay Fairness | Capacity-Demand Ratio | SL | UL | RL | RL Model | DL Model | ||
| [bh-dl] | FCNN | |||||||||
| [bh-dl2] | FCNN | |||||||||
| [bh-drl-d1] | DQN | FCNN | ||||||||
| [bh-drl-d2] | DQN | CNN | ||||||||
| [bh-drl-ga] | Proximal Policy Optimization | FCNN | ||||||||
| [bh-drl-mo1] | DQN | CNN | ||||||||
| [bh-drl-mo2] | DQN | CNN | ||||||||
| [bh-marl] | DDQN | FCNN | ||||||||
The key question of beam hopping is to find out which beams need to be activated when and for how long while maximizing the network performances given the capacity constraints [bh-ch]. This can be effectively formulated as an optimization problem considering different network performance metrics such as system throughput, delay, fairness, etc. as the objective(s) along with power and spectrum constraints. In [bh-op], a convex optimization framework with an objective to match the system capacity to traffic demand along with power allocation constraints is considered. This yields a close-form solution giving insights into resource allocation policies for maximizing network performances from different perspectives. However, from the perspective of real networks, the convex objective function is not very realistic, so the results are not applicable to real networks in a straightforward manner. Assuming the non-convexity of the problem, obtaining a globally optimal solution with efficient algorithms gets difficult. In [bh-gop], the steepest gradient descent algorithm is chosen to get the sub-optimal solution using the optimal set of precoding vectors. Some heuristic iterative approaches are also proposed in [bh-it1], [bh-it2], [bh-it3] to tackle these non-convex problems in a practical and feasible manner. Different meta-heuristic approaches like Genetic Algorithm (GA) [bh-ga], Simulated Annealing (SA) Algorithm [bh-sa], Particle Swarm Optimization (PSO) [bh-pso], and combined metaheuristic approaches like GA-SA [bh-gasa] have been considered to generate suboptimal solutions with a reduced amount of computational complexity.
@@
In [bh-dl], [bh-dl2], an SL approach is considered by forming labeled datasets with beam hopping patterns as outputs and channel matrix, transmission power, and traffic demand as inputs. First, a mixed integer linear problem formulation for matching the offered capacity to traffic demands is reduced to a simple linear programming problem. A training dataset is generated using conventional optimization algorithms and a DL model is trained on this dataset by considering beam hopping patterns as labels. Furthermore, the optimization framework can be potentially transformed into an RL problem to capture the optimal beam-hopping pattern in a learning environment. In [bh-drl-d1],[bh-drl-d2] the transmission delay is minimized considering the power and beam allocation constraints using a DRL approach. The state space consists of the average transmission delay and the buffer length with beam hopping pattern as actions and the negative Hadamard product of the current states, the negative of total queuing delay as the reward function. In [bh-drl-ga], a combined DRL-metaheuristic approach is considered to optimize both the throughput and delay fairness while at the same time designing different reward functions for the two cases. In [bh-drl-mo1], [bh-drl-mo2], a network consisting of real-time and non-real-time traffic is considered. A multi-objective problem minimizing the transmission delay for real-time traffics, maximizing the throughput for non-real-time traffics as well as overall delay fairness is considered. Individual reward functions are designed to capture each of the goals. In [bh-marl], a cooperative multi-agent framework is considered to dynamically allocate the power and bandwidth to illuminating beams optimizing throughput and delay fairness using a DDQN. In table V, we summarize the AI approaches for beam-hopping in NTN. As traffic demand changes with time, recursive architectures such as RNN, ESN, etc. should be also explored to design the NN for DL architectures used to address beam-hopping issues. Also, distributed learning architectures can be useful to design efficient beam-hopping schemes.
@@ V-C2 Spectrum Sharing
| Spectrum Sensing | Spectrum Occupancy Prediction | Spectrum Access | SL | UL | RL | RL Model | DL Model | ||
| [ntn-sp-dl] | CNN-BiLSTM | ||||||||
| [ntn-sr-dcnn] | CNN | ||||||||
| [ntn-ss-ccnls] | CNN-LSTM | ||||||||
| [ntn-spa-cnls] | CNN-LSTM | ||||||||
| [ntn-ss-drl1] | Modified Q-Learning | ||||||||
| [ntn-ss3] | SVM-CNN | ||||||||
| [ntn-ra-madrl2] | MADDPG | FCNN | |||||||
@@
In TNTN, the spectrum-sharing phenomenon needs more attention because we have a hierarchical network scenario consisting of non-terrestrial and terrestrial BSs as shown in Figure 3. To support this complex topology in a single framework, we need to come up with efficient spectrum-sharing strategies causing low interference to the users [ntn-ss-gen1]. In conventional spectrum sharing methods, we use efficient frequency reuse, leveraging directional antennas, adaptive power control, etc. methods to mitigate the effect of CCI. However, traditional four-color frequency reuse can effectively reduce the level of interference at the expense of more spectrum. beamforming approach can reduce the interference greatly, but that too comes at the cost of increasing complexity.
@@
To tackle this situation, a process called spectrum sensing is introduced in cognitive radio networks, where the unlicensed users can sense the occupancy status of the target band using some radio sensing method [cr-mit]. The popular spectrum sensing methods are Energy Detection (ED)[cr-ed], Cyclo-Stationary Detection (CSD) [cr-cyc], Eigen Value-based Detection (EVD) [cr-evd] etc. However, these methods either are simple with poor performance in low SNR scenarios (ED) or provide good performance but with more computational complexity (CSD and EVD). For these reasons, ML has been adopted for spectrum sharing to capture the correlation with a reduced computational complexity which can be extended to integrated satellite-terrestrial network scenarios [ntn-ss-gen1].
@@
Different intelligent learning approaches are adopted to tackle the spectrum-sharing problem for next-generation TNTN networks [ntn-ss1]. In [ntn-sp-dl], a spectrum-sharing strategy is developed for LEO satellites from the GEO satellite spectrum historical occupancy data using a CNN-BiLSTM model. Here the LEO satellite users are considered as unlicensed secondary users and the GEO satellite users are considered as the licensed primary users. In [ntn-ss-ccnls], a CNN-LSTM-based spectrum sensing method is introduced for satellites to capture the spatial and temporal correlation effectively for spectrum occupancy of satellite systems. In [ntn-spa-cnls], a CNN-LSTM model is introduced to predict the frequency assignment for satellites based on historical data. In [ntn-ss-drl1], a modified Q-Learning algorithm is used in an RL setup for the adaptive selection of access and modulation schemes for NGSO satellites in an NGSO-GEO system. In [ntn-ss3], an SVM model is first used for low complexity spectrum sensing, then a CNN-based spectrum prediction model based on historical data is developed. In [ntn-ra-madrl2], a cooperative MADRL framework is considered for bandwidth management in a game-theoretic model minimizing inter-beam interference. In [ntn-sr-dcnn], a CNN-based spectrum reconstruction method from incomplete data is discussed for satellite networks. In table VI, we summarize the AI approaches for spectrum sharing in TNTN. However, these spectrum sensing decisions need to be in real-time to increase the overall throughput of the secondary users; this means the conventional LSTM architectures need to be replaced by efficient low-complexity ESN architectures to tackle this in an online manner. Furthermore, spatial spectrum sharing scenarios need to be also considered along with the state of art temporal spectrum sharing scenarios leveraging the benefits of future 3D SAGIN networks.